OpenStudy (anonymous):

how to eliminate the arbitrary constant of this equation y= Ae^(2x) + Bcosx + Cxsinx to find the the d.e.

ganeshie8 (ganeshie8):

I think you need to differentiate it 3 times since you have 3 arbitrary constants

OpenStudy (michele_laino):

@hoho23 please try to calculate \[\frac{ d ^{2}y }{ dx ^{2}}\]

OpenStudy (anonymous):

y" = 4Aex^(2x) - B cosx - Cxsinx + 2Ccosx

OpenStudy (anonymous):

@eliassaab y" - 4y ?????

OpenStudy (anonymous):

@Michele_Laino then what's next??

OpenStudy (anonymous):

\[ y''(x)-4 y(x)=-5 \cos (x) (B+C x)-2 C \sin (x) \]

OpenStudy (michele_laino):

I'm trying to calculate the other higher derivatives of y

OpenStudy (anonymous):

You can choose A, B and C to be anything you want say A=B=C=1 and your DE will be \[ y''(x)-4 y(x)=-2 \sin (x)-5 (x+1) \cos (x) \]

OpenStudy (anonymous):

@eliassaab can i have a favor?? can you pls. show me the step by step how you come up with that answer?? plsssss :))

OpenStudy (anonymous):

i already computed the 5th derivative but i really don't know what to do next T.T

OpenStudy (michele_laino):

y''''-y=15Ae^(2x)-4Ccosx

OpenStudy (michele_laino):

and 2y''+2y=10Ae^(2x)+4C cosx

OpenStudy (anonymous):

You do not need to compute more than the second derivative. when you compute y''- 4 y you eliminate A and you can create many D.E that have y as a solution. As I said A, B and C can be any constants

OpenStudy (michele_laino):

@hoho23 I can write this: \[y ^{iv}+2y ^{ii}+y=25Ae ^{2x}\] may be useful?

OpenStudy (anonymous):

For example \[ y=e^{2 x}+x \cos (x)+\cos (x) \] is a solution of \[ y''-4 y=-5 \cos (x) (1+ x)-2 \sin (x) \]

OpenStudy (anonymous):

So there are infinite number of choices, unless you did not state the problem right.

OpenStudy (anonymous):

Does it say what order is the DE

OpenStudy (michele_laino):

y'(0)=2A

OpenStudy (anonymous):

@eliassaab no my professor said that we only need to eliminate the arbitrary content

ganeshie8 (ganeshie8):

I thought 3 arbitrary constants require a 3rd order DE ?

OpenStudy (anonymous):

constant rather

OpenStudy (anonymous):

You have to eliminate them to be a solution of a certain DE. Did he give you the DE?

OpenStudy (anonymous):

Otherwise, the problem has infinite number of solutions

ganeshie8 (ganeshie8):

i was thinking a second order DE would never require more than two arbitrary constants

OpenStudy (anonymous):

You see, when you solve a second order DE, you first solve the DE with the right hand side equal to zero and then you guess a special solution for the given right side.

ganeshie8 (ganeshie8):

yeah with a second degree eqn, wouldn't the DE satisfy only a specific set of values for A,B,C from the given general solution ? we can't call them arbitrary anymore right ?

ganeshie8 (ganeshie8):

im not so sure i understand the question clearly

ganeshie8 (ganeshie8):

*second order eqn

OpenStudy (anonymous):

how about this one y= Ae^(2x)cos3x + Be^(2x)sin3x

OpenStudy (anonymous):

@ganeshie8 pls help me with my follow up question :))

OpenStudy (michele_laino):

I think your ODE is: \[y ^{ii}-4y ^{i}+13y=0\]

OpenStudy (michele_laino):

@hoho23

OpenStudy (anonymous):

@Michele_Laino :))) thanks... michele what is your answer to my first question???

OpenStudy (michele_laino):

I tryed, nevertheless I don't found anything still

OpenStudy (michele_laino):

@hoho23 for your first question, i'm not sure, nevertheless try this: \[y ^{iii}-2(1+i)y ^{ii}-(1+4i)y ^{i}+2y=0\] where i is such that: \[i ^{2}=-1\] namely the immaginary unit

OpenStudy (michele_laino):

sorry the third coefficient is -(1-4i) and not -(1+4i)

OpenStudy (anonymous):

@Michele_Laino is that 3rd derivative or literally y^iii

OpenStudy (michele_laino):

it is the third derivative, namely \[\frac{ d ^{3}y }{ dx ^{3} }\]

OpenStudy (anonymous):

ah okay :))) thanks

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

@Michele_Laino thanks :)) hope for more help in the future :))

OpenStudy (michele_laino):

Ok!

OpenStudy (michele_laino):

@hoho23 Ok!, thanks!

OpenStudy (anonymous):

I think the DE has to have real coefficient.